OpenStudy (darkbluechocobo):

The distance that an object falls varies directly as the square of the time it has been falling. An object falls 65 feet in 2 seconds. Find how far it will fall in 5 seconds?

OpenStudy (darkbluechocobo):

406 feet 13 feet 25 feet 325 feet

OpenStudy (phi):

distance varies directly as t^2 means there is an equation d= k t^2 you need to first find "k" using the info 65 feet in 2 seconds then you can find d when t is 5

OpenStudy (darkbluechocobo):

65= k 4?

OpenStudy (darkbluechocobo):

oh wait it would be plugging in t for 5 so 65=k25

OpenStudy (phi):

yes, 65= k 4 divide both sides by 4 to get k= 65/4 now you know \[ d= \frac{65}{4} t^2 \] for t=5 , replace t with 5: \[ d= \frac{65}{4}5^2 \] and simplify

OpenStudy (darkbluechocobo):

65/4 = 16.25*25= 406.25 so 406

OpenStudy (phi):

yes, it looks like they rounded to the nearest foot.

OpenStudy (darkbluechocobo):

yes Well thank you for going through that :)

OpenStudy (phi):

sorry I lost track of this one, and forgot to get back to you.

OpenStudy (darkbluechocobo):

aha its fine :3